Suppose you’re evaluating three alternative MMMF investments. The first fund buys a diversified portfolio of municipal securities from across the country and yields 4.2 percent. The second fund buys only taxable, short-term commercial paper and yields 6.4 percent. The third fund specializes in the municipal debt from the state of New Jersey and yields 3.6 percent. If you are a New Jersey resident, your federal tax bracket is 36 percent, and your state tax bracket is 8 percent.

Calculate the sum of after-tax percentage yield of the three alternatives
Answer #1

Alternative 1: Diversified portfolio of municipal securities from across the country

In this case the federal tax of 36% won't be attracted. The investor will only be levied state tax at the rate of 8%

After-tax percentage yield = 4.2 - (4.2 * 8%) = 3.864%

Alternative 2: Taxable, short-term commercial paper

In this case both the federal tax of 36% and state tax of 8% will be levied. Thus total percentage of tax applicable will be 44% (36% + 8%).

After-tax percentage yield = 6.4 - (6.4 * 44%) = 3.584%

Alternative 3: municipal debt from the state of New Jersey

In this case no tax will be levied as it is tax exempt.

After-tax percentage yield = 3.60%

Sum of after-tax percentage yield of the three alternatives = 3.864% + 3.584% + 3.60% = 11.048%